# An introduction to infinite projected entangled-pair state methods for variational ground state simulations using automatic differentiation

### v3

@@ Appendix D Automatic differentiation in the language of differential geometry
$\begin{split}\int_{M}\bullet\leavevmode\nobreak\ \mu:&\leavevmode\nobreak\ \mathcal{F}(M)\xrightarrow{}\mathbb{R},\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \xmapsto{}\int_{M}f\mu,\end{split}$ (38)
@@
$v$ $=\sum_{i}v_{i}\frac{\partial}{\partial x_{i}}=\sum_{i}v_{i}e_{i},$ (40)
$df$ $=\sum_{i}\frac{\partial f}{\partial x_{i}}dx_{i}.$ (41)
@@
$dx_{j}\left(\frac{\partial}{\partial x_{i}}\right):=\frac{\partial x_{j}}{\partial x_{i}}=\delta_{i,j}.$ (42)
@@
$\bullet^{\flat}$ $:T_{p}M\xrightarrow{}T_{p}^{*}M,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ e_{i}\xmapsto{}dx_{i}=e_{i}^{\flat},$ (43)
$\bullet^{\sharp}$ $:T^{*}_{p}M\xrightarrow{}T_{p}M,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ dx_{i}\xmapsto{}e_{i}=dx_{i}^{\sharp}.$ (44)
@@
$\begin{split}\nabla f=\left(\sum_{i}\frac{\partial f}{\partial x_{i}}dx_{i}\right)^{\sharp}&=\sum_{i}\frac{\partial f}{\partial x_{i}}e_{i}\\ &=\left(\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{m}}\right),\end{split}$ (46)
@@
$(\underbrace{f_{*}\biggl{(}\overbrace{\frac{\partial}{\partial x_{i}}}^{\in T_{p}M}\biggr{)}}_{\in T_{f(p)}N})(g):=\frac{\partial}{\partial x_{i}}(f^{*}g)=\frac{\partial}{\partial x_{i}}(g\circ f).$ (48)
Now that we are equipped with the pullback and push-forward of differentials and derivatives we see how the gradient is calculated in the forward- and backward-mode AD. For this we will go back to our neat example from Sec. II.4 and slightly generalize. Say, we would like to take the gradient $\nabla E$ of a function that is composed of three primitive functions $E=f_{3}\circ f_{2}\circ f_{1}$. We say these primitive functions map between manifolds
$E:M_{1}\xmapsto{f_{1}}M_{2}\xmapsto{f_{2}}M_{3}\xmapsto{f_{3}}\mathbb{R}.$ (49)
@@
$f_{1}^{*}(f_{2}^{*}(df_{3}))=f_{1}^{*}((d(f_{3}\circ f_{2})))=d(f_{3}\circ f_{2}\circ f_{1})=dE.$ (51)
With our identification between tangent and cotangent vectors we finalize to $\nabla E=\left(dE\right)^{\sharp}$. If we express the differential that we start from $df_{3}$ in coordinates, we straightforwardly obtain the product of Jacobians as a result for the gradient. This also establishes the connection to the adjoint functions we talked about in the previous section and the vector-Jacobian product as discussed in Sec. II.4.
@@
$\begin{split}\sum_{i}f_{2*}\left(f_{1*}\left(\frac{\partial}{\partial x_{i}}\right)\right)(f_{3})\leavevmode\nobreak\ e_{i}&=\sum_{i}f_{1*}\left(\frac{\partial}{\partial x_{i}}\right)(f_{3}\circ f_{2})\leavevmode\nobreak\ e_{i}\\ &=\sum_{i}\frac{\partial}{\partial x_{i}}(\underbrace{f_{3}\circ f_{2}\circ f_{1}}_{=E})\leavevmode\nobreak\ e_{i}\\ &=\nabla E.\end{split}$ (53)

### v4

@@ Appendix D Automatic differentiation in the language of differential geometry
$\begin{split}\int_{M}\bullet\leavevmode\nobreak\ \mu:&\leavevmode\nobreak\ \mathcal{F}(M)\xrightarrow{}\mathbb{R},\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \xmapsto{}\int_{M}f\mu,\end{split}$ (D.2)
@@
$v$ $=\sum_{i}v_{i}\frac{\partial}{\partial x_{i}}=\sum_{i}v_{i}e_{i},$ (D.4)
$df$ $=\sum_{i}\frac{\partial f}{\partial x_{i}}dx_{i}.$ (D.5)
@@
$dx_{j}\left(\frac{\partial}{\partial x_{i}}\right):=\frac{\partial x_{j}}{\partial x_{i}}=\delta_{i,j}.$ (D.6)
@@
$\bullet^{\flat}$ $:T_{p}M\xrightarrow{}T_{p}^{*}M,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ e_{i}\xmapsto{}dx_{i}=e_{i}^{\flat},$ (D.7)
$\bullet^{\sharp}$ $:T^{*}_{p}M\xrightarrow{}T_{p}M,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ dx_{i}\xmapsto{}e_{i}=dx_{i}^{\sharp}.$ (D.8)
@@
$\begin{split}\nabla f=\left(\sum_{i}\frac{\partial f}{\partial x_{i}}dx_{i}\right)^{\sharp}&=\sum_{i}\frac{\partial f}{\partial x_{i}}e_{i}\\ &=\left(\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{m}}\right),\end{split}$ (D.10)
@@
$(\underbrace{f_{*}\biggl{(}\overbrace{\frac{\partial}{\partial x_{i}}}^{\in T_{p}M}\biggr{)}}_{\in T_{f(p)}N})(g):=\frac{\partial}{\partial x_{i}}(f^{*}g)=\frac{\partial}{\partial x_{i}}(g\circ f).$ (D.12)
Now that we are equipped with the pullback and push-forward of differentials and derivatives we see how the gradient is calculated in the forward- and backward-mode AD. For this we will go back to our neat example from Sec. 2.4 and slightly generalize. Say, we would like to take the gradient $\nabla E$ of a function that is composed of three primitive functions $E=f_{3}\circ f_{2}\circ f_{1}$. We say these primitive functions map between manifolds
$E:M_{1}\xmapsto{f_{1}}M_{2}\xmapsto{f_{2}}M_{3}\xmapsto{f_{3}}\mathbb{R}.$ (D.13)
@@
$f_{1}^{*}(f_{2}^{*}(df_{3}))=f_{1}^{*}((d(f_{3}\circ f_{2})))=d(f_{3}\circ f_{2}\circ f_{1})=dE.$ (D.15)
With our identification between tangent and cotangent vectors we finalize to $\nabla E=\left(dE\right)^{\sharp}$. If we express the differential that we start from $df_{3}$ in coordinates, we straightforwardly obtain the product of Jacobians as a result for the gradient. This also establishes the connection to the adjoint functions we talked about in the previous section and the vector-Jacobian product as discussed in Sec. 2.4.
@@
$\begin{split}\sum_{i}f_{2*}\left(f_{1*}\left(\frac{\partial}{\partial x_{i}}\right)\right)(f_{3})\leavevmode\nobreak\ e_{i}&=\sum_{i}f_{1*}\left(\frac{\partial}{\partial x_{i}}\right)(f_{3}\circ f_{2})\leavevmode\nobreak\ e_{i}\\ &=\sum_{i}\frac{\partial}{\partial x_{i}}(\underbrace{f_{3}\circ f_{2}\circ f_{1}}_{=E})\leavevmode\nobreak\ e_{i}\\ &=\nabla E.\end{split}$ (D.17)